40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 307     Entry time: Thu Aug 26 19:58:06 2010
Author: Frank 
Type: Notes 
Category: BEAT 
Subject: changed ACAV setpoint temperature - both cavities locked 

changed the ACAV temp to 37.3 degrees. Both cavities are locked now.

I'll monitor the VCO input signal to center the range by adjusting the temp a little bit more.

ELOG V3.1.3-